DIRC Operations Weekly Report

Period: Wednesday 2007/1/10 - Wednesday 2007/1/17

Oncall: Jose Benitez
Fast Monitoring: Jose Benitez, Alejandro Perez

File location: $BFROOT/www/Detector/DIRC/WeeklyReports/report_20070117.html
Plot directory: $BFROOT/www/Detector/DIRC/WeeklyReports/Plots/report_20070117 (empty)

Contents


BaBar-PEP news

PEP peak lumi: 0 x 1033 (previous week: NA)

BaBar integrated lumi: 0 fb-1 (previous week: NA)

Run 6 PEP delivery: 0 fb-1

Run 6 total delivery: 0 fb-1 Run 6 total BaBar recorded: 0 fb-1 including 0 fb-1 off-peak data

PEP activities this week:

  • Run 6 startup tuning. Beams soon.

    Foreseen work opportunities (PEP down periods):

  • Day shift Jan 17 & 31.

    BaBar activities/issues this week:

  • DIRC HV half way (V1) since this weekend while PEP is trying to inject.

    DIRC activities

    See DIRC logbook for details

    Backgrounds

    NA

    Scheduled activities

    Pending Work

    Manpower situation

    Bradley, Jose and Alejandro rotating backup & on-call commissioning duties, with Alejandro working doubletime.

    Links